# Preparing for Super-Reactivity: Early Fault-Detection in the Development of Exceedingly Complex Reactive Systems

### v1

@@ Abstract
Super-Reactive Systems, System Engineering, Software Engineering, LLM, Models, Specifications, Interaction, Simulation, Verification, Autonomous systems, Systems of Systems.
@@ I Introduction
Since the 1985 identification of the category of reactive systems[1], a plethora of methods, languages and tools have been introduced to support the development of such systems. Today, complex reactive systems are penetrating almost every aspect of life, including communications, commerce, finance, healthcare, aviation, land transportation, manufacturing, entertainment, and more. The complexity of new systems is compounded by the fact that they are interwoven with other systems and with the physical world, and are constantly changing and evolving. We term this emerging kind of system super-reactive (SR). While system and software engineering (SySE) is benefitting from new developments in generative AI and large language models (LLMs), the challenge of building safe and reliable SR systems remains open. Despite applying the best tools and methodologies, any given system is likely to conceal undesired and very often even unsafe behaviors and impending failures, with the risk of adverse effects on human life, the environment, society and the economy. Thus, while early discovery and handling of such faults is required, it remains a tantalizing challenge. In this paper, we propose tackling this issue relying on the following principles: (1) Allow model elements that are expressed in natural language (NL), benefitting from the expressive power of natural language, its sensitivity to delicate context variations and its ability to navigate multiple levels of abstraction, and carrying out just-in-time (JIT), deferred, interpretation of such NL elements. (2) Discover and document otherwise-tacit interdependencies among separately specified, seemingly orthogonal requirements.
@@ II The Problem
Current AI-based solutions assist in various activities of development, including code generation to debugging, with a prominent example being the GitHub Copilot; see, also, e.g., [21] and references therein. However, applying such tools in the context of early specification is mostly limited to automated modeling, which we discuss in Section III-C.
@@ III The roadmap
Below we list the elements of an approach and an architecture for modeling SR systems, including a set of intelligent tools for simulation and analysis, which, together, can enable the much desired early preemptive discovery of hidden faults, while coping with the existing challenges and technology gaps. In bounding the scope of the proposed solution, we exclude the use of AI, ML and LLMs in runtime decision making, monitoring, DevOps, or the formal verification of final code . And, while some reasoning functions of the proposed solutions may be similar to those carried out by expert human engineers and domain professionals, we focus on enabling presently impractical or impossible analyses, and much less on automation of manual tasks.
@@ III-A The Intelligent Development Aide
The Intelligent Development Aide (IDA) is a shared layer of services that offers the following to the overall solution: (i) intelligence, including learning, inference, and generative abilities; (ii) NL-based interaction capabilities; and, (iii) general world knowledge and certain domain-specific expertise. The IDA will rely on present and future technologies and tools that come under the umbrella of AI, Generative AI, Machine Learning, Deep Learning, Large Language Models (LLMs), etc. It will be constructed, among other things, by fine-tuning, enhancing and extending AI-based tools, relying on techniques like those of  [24, 25, 26, 27, 28] and future emerging ones. With inputs coming from diverse systems, with textual and visual depictions of execution scenarios and logs of normal and faulty interactions, the IDA will be trained to recognize unique software and system engineering issues and new delicate kinds of interdependencies.
@@ III-B The i-model
And third, modularity, encapsulation, and logical decomposition are central principles in software engineering, and in engineering in general. However, separately specified requirements often have tacit, unstated dependencies, which show up as exceptions, priorities, alternatives, complementary or concurrent actions, mutually exclusive conditions, etc. It was up to the engineers to infer these implicit relations, and to reflect their understanding in the implementation. In our automated construction of i-models from a wide range of specifications, special focus will be put on discovering such unstated relationships and capturing them in the model, despite the entanglement that they may imply.
The i-models will store diverse information, including requirements, goals, behaviors, scenarios, properties, emergent effects, etc., as well as groupings, abstractions, and relations of such entities. It will also contain meta information about potential changes due to evolution of the system and its environment, allowing further analysis of potential future trajectories.
@@
The management of the immense knowledge stored in the i-model will be divided among three realms: (i) the entities themselves, including structured data and unstructured documents, much of which can be in NL; (ii) the rich relationship between entities, which will be represented in the i-model database; and, (iii) the general and application-specific knowledge captured in the IDA components, both in advance, and following the training associated with the building and analysis of the i-model.
@@ III-C The i-model builder
To illustrate related capabilities of AI tools we conducted several experiments. For example, we described to an LLM requirements and properties for the edges and nodes of graphs. The LLM was able to produce the full list of nodes and edges; its few mistakes highlighted omissions and ambiguity in the specification. For example, we prompted the LLM as follows: “… Consider the set S containing the entities A, B, C, D. Consider all subsets A1-A15 of S containing at least one element. … Consider the following graph. From each subset A_i there is an undirected edge to each subset that has fewer entities than A_i (they do not have to be A_i’s subsets). [Furthermore], all the subsets with exactly 2 elements are connected to each other with directed edges in a cycle - i.e., each one has exactly one entering arrow and one exiting arrow. The order does not matter…”.
The LLM’s response consisted of a complete list of nodes, followed by the list of edges, including, for example, {A,B}: Directed edge to {A,C}; {A,C}: Directed edge to {A,D}”. We then asked it to describe a higraph [34] having four basic nodes, blobs around certain sets of nodes, and certain edges between blobs. To make the problem real, we cast the blobs as assignments of persons to projects, and the edges as dependencies among projects. The LLM was able to create a bulleted list of all the blobs and their contents, and connecting edges.
We expect the construction of the i-model builder to extend the growing amount of research on using GAI and LLMs for creating models from NL inputs for; see, e.g.  [35, 28, 36, 37, 38, 39]
@@ III-D The i-model execution and simulation engine
This component carries out dynamic, deferred, just-in-time (JIT) interpretation of the definite and the NL information in the i-model, and outputs descriptions of possible runs and the unfolding of “what if” scenarios at desired abstraction levels, in the forms of narratives, event logs, and even synthetic videos.
@@
For illustration, consider the following example, which we ran on OpenAI ChatGPT. Prompt (slightly edited): “Consider a 3 lane street. 30 equal-size passenger cars are filling the street, standing in front of a traffic light. No cars arrive from the back of the queue throughout this scenario. The third car stalls and cannot move. No one stops to help. Describe the behaviors and positions of the 30 cars after the light turns green. Assume the light is green for a very long time.”.
Here are a few excerpts from the LLM’s response: “When the traffic light turns green: …
@@
1. Cars in the Middle Lane (Lane 2): (a) The first two cars in the middle lane (Cars 1 and 2) will start moving forward as the light turns green. (b) The third car in the middle lane (Car 3) will not move… (c) The cars behind the third car … will initially wait, expecting the third car to move. After realizing that it is not moving, they will either: (i) Start honking, …; (ii) Some drivers may try to change lanes … (iii) Some drivers may remain stuck behind the third car if changing lanes is not possible…
2. Cars in the Left Lane… and Right Lane…: (a) The cars in these lanes … will begin moving forward … (b) The first few cars … positioned parallel to the stationary third car, will drive past it without much issue. (c) The cars behind them will continue moving forward, potentially picking up speed…”
@@
When we divided the road into small virtual cells, the LLM was able to describe car movements through the numbered cells, making the textual scenario more concrete and mathematically analyzable.
@@
The LLM responses also contained errors; for example, the LLM suggested the possibility that cars blocked behind the stalled car will not be able to change lanes at all, ignoring the specified facts that the traffic light stays green for a long time, and that no new cars arrive during the scenario in question; however, we believe that with proper training, responses can be made highly reliable. And in any case, the answers can be checked and validated interactively, and domain-specific knowledge can be added. Note how the LLM incorporated into the responses its domain knowledge; for example, the adding the possibilities of changing lanes and honking, which were not included in the original specification.
The execution engine will, of course, benefit from state-of-the-art execution and simulation techniques, like those in SysML, Statecharts, and Scenario-based programming[40, 41], as well as in direct execution of NL specifications as in [27] and references therein.
@@
Borrowing from techniques for test-case generation [42], execution and simulation engine will also generate and process batches of diverse, yet relevant, “what if” scenarios, and store their execution results for further processing.
@@ III-E The i-model analysis engine
In our experiments, an LLM was able to discover paths and absence thereof in graphs like those built in the example of Section III-C. Other analytic LLM capabilities that can support i-model analysis are described in papers like [37]. These include computing when two independent periodic events may occur simultaneously, explaining behavior, articulating system properties, etc. Furthermore, it is expected that LLM general and domain-specific analytical capabilities will be extended and deepened, and are likely to be intertwined with ongoing research in software and system engineering. Developments along the present roadmap can incorporate such enhancements, and target them specifically at early fault-detection.
@@ IV Conclusion
We are now in the process of initiating a research project following the above roadmap. Development of models and tools that enable simulation and analysis of highly complex systems based only on early specifications can dramatically enhance our ability to develop reliable, safe, and productive SR systems. A combination of the recent advances in AI and a fresh perspective on what may and may not qualify as a model entity or be acceptable as a simulation result, may enable the achievement of this tantalizing goal.

### v2

@@ Abstract
System Engineering, Software Engineering, LLM, Models, Specifications, Interaction, Simulation, Verification, Autonomous systems, Systems of Systems.
@@ I Introduction
Since the 1985 identification of the category of reactive systems[1], a plethora of methods, languages and tools have been introduced to support the development of such systems. Today, complex reactive systems are penetrating almost every aspect of life, including communications, commerce, finance, healthcare, aviation, land transportation, manufacturing, and more. The complexity of new systems is compounded by the fact that they are interwoven with other systems and with the physical world, and are constantly changing and evolving. We term this emerging kind of system super-reactive (SR). While system and software engineering (SySE) is benefitting from new developments in generative AI and large language models (LLMs), the challenge of building safe and reliable SR systems remains open. Despite applying the best tools and methodologies, any given system is likely to conceal undesired and very often even unsafe behaviors and impending failures, with the risk of adverse effects on human life, the environment, society and the economy. Thus, while early discovery and handling of such faults is required, it remains a tantalizing challenge. In this paper, we propose tackling this issue relying on the following principles: (1) Allow model elements that are expressed in natural language (NL), benefitting from the expressive power of natural language, its sensitivity to delicate context variations and its ability to navigate multiple levels of abstraction, and carrying out just-in-time (JIT), deferred, interpretation of such NL elements. (2) Discover and document otherwise-tacit interdependencies among separately specified, seemingly orthogonal requirements.
@@ II The Problem
Current AI-based solutions assist in various activities of development, including code generation to debugging, with a prominent example being the GitHub Copilot; see, also, e.g., [21] and references therein. However, applying such tools in the context of early specification is mostly limited to automated modeling, discussed in Section III-C.
@@ III The roadmap
Below we list the elements of an approach and an architecture for modeling SR systems, including a set of intelligent tools for simulation and analysis, which, together, can enable the much desired early preemptive discovery of hidden faults, while coping with the existing challenges and technology gaps.
In bounding the scope of the proposed solution, we exclude the use of AI, ML and LLMs in runtime decision making, monitoring, development operations (DevOps), or the formal verification of final code. And, while some reasoning functions of the proposed solutions may be similar to those carried out by expert human engineers and domain professionals, we focus on enabling presently impractical or impossible analyses, and much less on automation of manual tasks.
@@ III-A The Intelligent Development Aide
The Intelligent Development Aide (IDA) is a shared layer of services that offers the following to the overall solution: (i) intelligence, including learning, inference, and generative abilities; (ii) NL-based interaction; and, (iii) general world knowledge and certain domain-specific expertise. The IDA will rely on present and future technologies that come under the umbrella of AI, Generative AI, Machine Learning, Deep Learning, Large Language Models (LLMs), etc. It will be constructed, among other things, by fine-tuning, enhancing and extending AI-based tools, relying on techniques like those of  [24, 25, 26, 27, 28] and future emerging ones. With inputs from specifications of diverse systems, with textual and visual depictions of normal and faulty execution scenarios, the IDA will be trained to recognize unique software and system engineering issues and new delicate kinds of interdependencies.
@@ III-B The i-model
And third, modularity, encapsulation, and logical decomposition are central principles in software engineering, and in engineering in general. However, separately specified requirements often have tacit, unstated dependencies, which show up as exceptions, priorities, alternatives, complementary or concurrent actions, mutually exclusive conditions, etc. It is commonly up to the engineers to infer these implicit relations, and to reflect their understanding in the implementation. In our automated construction of i-models from a wide range of specifications, special focus will be put on discovering such unstated relationships and capturing them in the model, despite the entanglement that they may imply.
The i-models will store diverse information, including requirements, goals, behaviors, scenarios, emergent properties, etc., as well as groupings, abstractions, and relations of such entities. It will also contain meta information about potential changes due to evolution of the system and its environment, allowing further analysis of potential future trajectories.
@@
The immense knowledge stored in the i-model will be divided among three realms: (i) the entities themselves, including structured data and unstructured NL documents; (ii) the relationships between entities, represented in the i-model database; and, (iii) the general and application-specific knowledge captured in the IDA components, both in advance, and following the building and analysis of a given i-model.
@@ III-C The i-model builder
We expect the construction of the i-model builder to extend the growing amount of research on using GAI and LLMs for creating models from NL inputs; see, e.g.  [34, 28, 35, 36, 37, 38].
@@ III-D The i-model execution and simulation engine
This component performs dynamic, deferred, just-in-time (JIT) interpretation of the definite and the NL information in the i-model, and outputs descriptions of possible runs and the unfolding of “what if” scenarios at desired abstraction levels, as narratives, event logs, and even synthetic videos.
@@
For illustration, consider the following example, which we ran on OpenAI ChatGPT. Prompt (slightly edited): “Consider a 3 lane street. 30 equal-size passenger cars are filling the street, standing in front of a traffic light. No cars arrive from the back of the queue throughout this scenario. The third car stalls and cannot move. No one stops to help. Describe the behaviors and positions of the 30 cars after the light turns green. Assume the light is green for a very long time.”
@@
Here are a few excerpts from the LLM’s response: “When the traffic light turns green:…1. Cars in the Middle Lane…: (a) The first two cars in the middle lane (Cars 1 and 2) will start moving forward as the light turns green. (b) The third car in the middle lane (Car 3) will not move…(c) The cars behind the third car…will initially wait, expecting the third car to move. After realizing that it is not moving, they will either: (i) Start honking,…; (ii) Some drivers may try to change lanes…(iii) Some drivers may remain stuck behind the third car if changing lanes is not possible… 2. Cars in the Left Lane…and Right Lane…: (a) The cars in these lanes…will begin moving forward… (b) The first few cars…positioned parallel to the stationary third car, will drive past it without much issue. (c) The cars behind them will continue moving forward, potentially picking up speed…”
@@
When we divided the road into small virtual cells, the LLM was able to describe car movements through the numbered cells, making the scenario more mathematically analyzable.
@@
The LLM responses also contained errors; for example, they suggested that cars blocked behind the stalled car may not be able to change lanes at all, ignoring the specification that the traffic light stays green indefinitely, and no new cars arrive during the scenario; however, we believe that with proper training, responses can be made highly reliable. And in any case, the answers can be checked and validated interactively, and domain-specific knowledge can be added. Note how the LLM incorporated into the responses its domain knowledge; for example, adding the possibilities of changing lanes and honking, which were not in the original specification.
@@
The execution engine will, of course, benefit from state-of-the-art execution and simulation techniques, like those in SysML, Statecharts, and Scenario-based programming[39, 40], or in direct execution of NL specifications as in [27] and references therein. Borrowing from techniques for test-case generation [41], the execution engine will also generate and process batches of diverse, yet relevant, “what if” scenarios, and store their execution results for further processing.
@@ III-E The i-model analysis engine
For example, we described to ChatGPT two parallel synchronous state machines. With some trial and error, and with checking and corrections by engineers, the LLM was able to answer whether certain composite states were reachable or not, describe relevant paths, and construct the full state graph of the composite machine with all composite states and transitions. Other analytic LLM capabilities that can support i-model analysis are described in papers like [36, 42]. These include computing when two independent periodic events may occur simultaneously, explaining behavior, articulating system properties, checking model consistency, etc. Furthermore, it is expected that LLM general and domain-specific analytical capabilities will be extended and deepened, and they are likely to be intertwined with ongoing research in software and system engineering. Developments along the present roadmap can incorporate such enhancements, and target them specifically at early fault-detection.
Still, automated validation techniques must be researched and developed. That is, assume that the fault-detection solution is confirmed to work well on reasonably tractable models, like compositions of small specifications. Can one trust the solution’s answers on larger problems? And wouldn’t a trusted automated external validation tool make the AI-based solution unnecessary altogether? We believe that with a combination of well-documented abstraction relations, AI-explainability, randomized testing of model answers, and powerful projection of relevant model perspectives, one can create high confidence in the model’s answers.
@@ IV Conclusion
We are now in the process of initiating a research project following the above roadmap. Development of models and tools that enable simulation and analysis of highly complex systems based only on early specifications can dramatically enhance our ability to develop reliable, safe, and productive super-reactive systems. A combination of the recent advances in AI and a fresh perspective on what may and may not qualify as a model entity or be acceptable as a simulation result, may enable the achievement of this tantalizing goal.